# EFSA: Episodic Few-Shot Adaptation for Text-to-Image Retrieval

### v1

@@ 1 Introduction
A key observation about models like CLIP and SigLIP [41] is that they excel at distinguishing and identifying easy negatives—dissimilar images for text-to-image retrieval tasks, as reflected in their strong performance for higher recall metrics such as Recall@5 and Recall@10 as can be seen from Figure 1. However, they often struggle to correctly rank hard negatives, which are images with subtle similarities or minor variations compared to the query, affecting Recall@1. This limitation becomes especially problematic in an open-domain setting where visually similar images across domains may vary slightly, thus requiring more fine-grained alignment.
@@
  • We identify limitations in current text-to-image retrieval benchmarks, emphasizing the need for evaluation within a realistic, open-domain setting that encompasses diverse and complex visual content.
  • We introduce Episodic Few-Shot Adaptation (EFSA), a novel test-time adaptation framework which enhances the model robustness against hard negatives, by episodically learning from them.
@@ 2.1 Text-to-image retrieval
Although these datasets are valuable for evaluating model capabilities in natural image settings, their limitations in terms of size and coverage hinder a comprehensive evaluation of model performance across diverse domains. Our work aims to address this gap by expanding the scope of retrieval evaluation across multiple domains, thereby assessing model performance in more realistic and challenging open-domain scenarios.
@@ 2.2 Adaptation methods
Continual learning also addresses domain shifts and multi-domain scenarios, by incrementally tuning a model to new domains to mitigate catastrophic forgetting [45]. Yet, this approach assumes knowledge of the relevant domains and access to training data, both impractical for T2I retrieval in real-world scenarios.
@@ 2.3 Synthetic captions for retrieval
Various recent works have shown that continual pre-training of VLMs like CLIP on synthetically generated captions yields improved performance on a range of tasks, including retrieval [4, 42, 37]. Using a generative VLM like LLaVa [22], these works re-label millions of images with well-formed detail captions, far more informative than the noisy captions used for the initial pre-training of CLIP, for example [28]. While continued pre-training allows the modified VLM to perform better on average across many tasks and domains, it does not guarantee optimal performance in any one of these tasks and domains. Our episodic few-shot training framework leverages just a few highly relevant synthetically captioned images to tune the VLM to the specific domain of the query.
@@ 3.1 Preliminaries
Contrastive Language-Image Pre-training (CLIP) comprises two encoders: the visual encoder $\mathcal{F}_{\theta_{v}}$, which maps visual input $X$ to a fixed-length representation $\bm{f}_{v}$, and the text encoder $\mathcal{F}_{\theta_{t}}$, which processes text input, $T$, and generates a latent textual feature $\bm{f}_{t}$. In zero-shot text-to-image retrieval, a string of text, $q$, is used to query a pool of images, $\mathcal{I}$. Given the representations of the query, $\bm{f}_{q}$, and of all images $\{\bm{f}_{v,i}\}_{i\in\mathcal{I}}$, a cosine similarity score $s_{i}=\mathtt{sim}(\bm{f}_{q},\bm{f}_{v,i})$ is computed for each image, $i$, to form a final ranking.
@@ 3.2 EFSA: Episodic Few-Shot Adaptation
Although CLIP-style models achieve impressive zero-shot performance across varied downstream tasks, they struggle to accurately rank hard negative pairs in text-to-image retrieval settings, particularly in complex, multi-domain scenarios. We address this limitation with EFSA, an adaptation framework which utilizes top-$k$ highly similar images as hard negatives, to improve multi-modal alignment in the specific domain of the query. In this section, we detail our three-stage EFSA methodology, explaining how each stage contributes to more robust retrieval across diverse and challenging domains. For a visualization of the approach, see Figure 2.
Initial Top-$k$ Retrieval In the initial zero-shot retrieval step, we compute the score, $s$, between a given query text, ${q}$, and all images $\bm{i}\in\mathcal{I}$. The top-$k$ most similar images are selected, forming the set $\mathcal{I}_{\text{top}}=\{\bm{i}_{1},\bm{i}_{2},\dots,\bm{i}_{k}\}$. Given a large-enough $K$ and a sufficiently generalized VLM, we can expect the ground truth to be in $\mathcal{I}_{\text{top}}$, but possibly not at the top rank, where it should be. Crucially, the rest of the images in the set would be similar to each other and to the ground truth, thus forming a set of domain-specific hard negatives with respect to the ground truth. In the next step, we leverage these hard negatives to adapt the VLM to the relevant visual domain.
Data Augmentation and Episodic Training In this step, pseudo-captions $\mathcal{C}_{\text{top}}=\{\bm{c}_{1},\bm{c}_{2},\dots,\bm{c}_{k}\}$ are generated for each image in $\mathcal{I}_{\text{top}}$ using a pre-trained image captioning model, guided by a predefined prompt $\mathcal{P}$, chosen to ensure relevant caption generation. Together $\mathcal{I}_{\text{top}}$ and $\mathcal{C}_{\text{top}}$ form a dataset of aligned image-caption pairs, which are used to adapt the VLM with a two-fold objective, using contrastive and hinge losses.
@@
where $\text{sim}(\cdot,\cdot)$ denotes cosine similarity between the visual feature $\bm{f}_{v,i}$ and text feature $\bm{f}_{t,i}$, and $\tau$ is the temperature parameter.
@@
The final test-time adaptation loss, $\mathcal{L}_{\text{test}}$, is a weighted combination of these losses:
$\mathcal{L}_{\text{test}}=\alpha\,\mathcal{L}_{\text{contrastive}}+\beta\,\mathcal{L}_{\text{hinge}},$ (3)
where $\alpha$ and $\beta$ balance the contribution of each component.
These objectives are used to learn a set of LoRA layers [14] in both $\mathcal{F}_{\theta_{v}}$ and $\mathcal{F}_{\theta_{t}}$ in the VLM. With LoRA, the VLMs is adapted to the specific domain of the query at a low computational cost, with reduced risk of overfitting to noise in $\mathcal{I}_{\text{top}}$ and $\mathcal{C}_{\text{top}}$.
Inference In this final step, the updated VLM is used to re-rank the images in $\mathcal{I}_{\text{top}}$ with respect to the query $q$, leveraging domain-adapted representations to enhance retrieval recall at the top ranks.
Once inference is complete for a particular query, the LoRA weights of visual and text encoder are reset, ensuring that subsequent queries start from a neutral state, allowing the model to adapt to each new query independently, without cross-domain interference.
@@ 4.1 Experimental Setup
Implementation Details We use CLIP-B16 as the main VLM in all experiments, but additionally test the generalization of our approach to SigLIP [41] as well (see Supplementary §8). LoRA layers are added to all attention layers and multi-layer perceptron components, in both the vision and text encoders. The LoRA parameters are initialized with Xavier initialization, using a rank of $r=64$ and a scaling factor of 15. The loss function is updated in a single step with parameters $\alpha=1.7$ and $\beta=0.3$, using the AdamW optimizer and a learning rate of $5\times 10^{-4}$. All hyperparameters were tuned on the validation splits of COCO and Flickr30k, adopting the values that performed best across the two datasets on average.
For top-$k$ sampling, we set $k=16$. Pseudo-captions are generated with the LLaVA 1.5-13B model [22], using the prompt: ”Describe what you see in detail with a maximum of 30 words.” As the captions are independent of the queries, they can be generated offline and cached. Similarly, we precompute and cache the representations of all images in the retrieval pool, as every initial top-$k$ retrieval step relies on the same pre-trained VLM representations.
@@
Experimental Settings We include two experimental settings: a single-domain setting, in which the image pool contains only images from a single, domain-specific dataset, allowing us to assess performance within a focused domain; and a more challenging and realistic multi-domain setting, in which we expand the image pool to include images from all datasets listed in Table 1.
@@
Below, we compare our approach to these baselines using the standard retrieval metric Recall@$k$ for $k=1,5,10$, with special attention to Recall@1, since EFSA focuses on re-ranking the top retrieval candidates for an optimal prediction in the top-$1$ position.
@@ 4.2 Results
The main results are shown in Table 2 for the single-domain setting, which treats each dataset independently, and Table 3 for the multi-domain setting, where the retrieval pool spans multiple diverse datasets.
Single-Domain Setting While this setting is not the ultimate target of our work, it lays the foundation for the follow-up discussion on multi-domain results. The first thing to note here is the considerable variation in zero-shot scores across the different datasets. The Books, ArtCap and SciCap datasets, in particular, prove far more challenging than COCO and Flickr30k, underscoring the importance of domain diversity in T2I retrieval evaluation.
@@
RLCF delivers improvements over the zero-shot baseline for 6 out of 8 datasets, but these improvements are mostly negligible in comparison to those seen with the fine-tuning baseline. The one exception beign SciCap where RLCF indeed proves best overall, yielding a Recall@1 improvement of 6.54 points. Meanwhile, however, for NASA and TextCap, RLCF exhibits degenerate behavior, scores plummeting for these datasets across all Recall@k metrics. This instability in the reinforcement-based optimizaiton of RLCF leaves simple fine-tuning as the strongest baseline on average for us to compare against.
@@
In the multi-domain setting, EFSA proves best on 6 out of 8 datasets, with an average Recall@1 4.27 points over the zero-shot baseline (27.56 v. 23.29). This number coincidentally matches exactly the improvement of EFSA over the zeros-shot baseline in the single-domain setting. In light of the absolute drop in scores in the multi-domain setting compared to the single-domain setting, this stable improvement with EFSA indicates that our method is highly robust to noise in the $\mathcal{I}_{top}$ subset: even if some lower-quality candidates get retrieved, EFSA successfully adapts to the domain-specific patterns and ranks better candidates higher.
@@
Figure 3: Qualitative comparison of EFSA and zero-shot CLIP. Green-framed images are the ground-truths for each text query (shown on top). Our method successfully re-ranks ground-truth images to the first rank, outperforming zero-shot CLIP.
@@ 4.3 Qualitative Analysis
Figure 3 provides an insight into the performance gains achieved with EFSA. Looking at the top-4 predictions retrieved with the zero-shot method and with EFSA for two text queries from the Flickr30k dataset, we see that EFSA picks up on fine detail such as the presence of a glass of wine in the top example, and the orientation of the man in the second example.
@@ 5 Ablations
We perform extensive empirical analysis and ablation studies to evaluate how various design choices influence our method’s performance. We perform the ablation on a subset of domains: COCO, NASA, SciCap, and ArtCap.
@@
Figure 4: Comparison of LoRA parameter tuning versus tuning all model parameters across multiple datasets demonstrates that full parameter tuning on limited data compromises the model’s generalizability and performance.
5.2 Effect of Loss Function
Loss Function COCO ArtCap
R@1 R@5 R@10 R@1 R@5 R@10
Hinge 30.15 50.78 57.79 10.91 22.73 27.47
Contrastive 28.23 49.20 56.47 10.33 21.66 26.61
Combined 30.14 50.96 57.82 11.13 22.83 27.50
Table 5: Comparison of recall performance at incremental top-$k$ values across datasets. We find that recall improves as the value of $k$ increases, particularly for higher ranks.
@@ 5.3 Top-k Selection
Retrieval performance measured in terms of Recall@1 exhibits a more nuanced pattern. While Recall@1 scores initially improve as the top-$k$ value increases, the gains plateau or even decline beyond a certain top-$k$. For example, on COCO, Recall@1 reaches a peak at $k=16$ with a score of 30.14 but drops slightly as $k$ continues to increase. This suggests that while expanding the retrieval pool improves general recall, very large top-$k$ values may introduce additional noise, compromising precision at the top rank. Thus, selecting an optimal top-$k$ is essential to balance high precision among top-ranked candidates, with broader recall across the retrieval set.
@@ 5.4 Effect of Epoch
All experiments so far were performing with a single epoch of training. In Figure 5 we explore whether increasing the number of epochs has a positive impact on performance, with the finding that by and large that is not the case. For two datasets, COCO and ArtCap, we see slightly higher scores at Recall@1 and Recall@5 on the second epoch of training, but the general trend is for recall to drop with more extensive training. Interestingly, the drop is more pronounced at higher ranks. In NASA, for example, Recall@1 drops by less than 10 points across the four epochs of training, while Recall@5 drops by over 25 points. This indicates that prolonged training on a limited or domain-specific set of images can cause the model to memorize specific features rather than develop more generalized representations, robust across various types of images. Regardless, from a computational point of view, having to perform a single epoch of training to reap the benefits of EFSA, is optimal.
6 Limitations
Our approach relies on an external image captioning model to generate synthetic captions. The quality of these captions may impact the performance of EFSA. Notably, we employ an open-source model (LLaVA), which proves effective in this context and which future work can also rely on.
Like any test-time adaptation method, our approach updates model weights for each test sample to achieve optimal performance, thus incurring a higher computational cost compared to zero-shot inference. Notice that compared to RLCF, for example, the exact same computational budget is needed for the forward passes (given a fixed top-$k$), while our method is more computationally efficient in the backward pass, as we only update a fraction of the model parameters in the form of LoRA layers, rather than the entire vision encoder, as done in RLCF.

### v3

@@ 1 Introduction
A key observation about models like CLIP and SigLIP [41] is that they excel at distinguishing and identifying easy negatives—dissimilar images for text-to-image retrieval tasks, as reflected in their strong performance for higher recall metrics such as Recall@5 and Recall@10 (see Figure 1 for results on three distinct datasets). However, they often struggle to correctly rank hard negatives, which are images with subtle similarities or minor variations compared to the query, affecting Recall@1. This limitation becomes especially problematic in an open-domain setting where visually similar images across domains may vary slightly, thus requiring more fine-grained alignment.
@@
  • We identify limitations in current text-to-image retrieval benchmarks, emphasizing the need for evaluation in a realistic, open-domain setting of diverse visual content.
  • We introduce EFSA, a novel test-time adaptation framework which enhances model robustness against hard negatives, by episodically learning from them.
@@ 2.1 Text-to-Image Retrieval
Although these datasets are valuable for evaluating model capabilities in the natural image domain, their limitations in terms of size and coverage hinder a comprehensive evaluation of model performance. Our work addresses this gap by expanding the scope of retrieval evaluation to multiple domains, to assess model performance in a more realistic and challenging open-domain scenario.
@@ 2.2 Adaptation Methods
Continual learning also addresses multi-domain scenarios, by incrementally tuning a model to new domains to mitigate catastrophic forgetting [46]. Yet, this approach assumes knowledge of the relevant domains and access to training data, both impractical for real-world T2I retrieval.
@@ 2.3 Synthetic Captions for Retrieval
Recent work has shown that continual pre-training of VLMs like CLIP on synthetically generated captions yields improved performance on a range of tasks, including retrieval [4, 42, 37]. Using a generative VLM like LLaVa [22], these works re-label millions of images with well-formed detailed captions, far more informative than the noisy captions used for the initial pre-training of CLIP [28]. While continued pre-training allows the modified VLM to perform better on average across many tasks and domains, it does not guarantee optimal performance in any one of these tasks and domains. Our episodic few-shot training framework leverages just a few highly relevant synthetically captioned images to tune the VLM to the specific domain of the query.
@@ 3.1 Preliminaries
Contrastive Language-Image Pre-training (CLIP) comprises two encoders: the visual encoder $\mathcal{F}_{\theta_{v}}$, which maps visual input $X$ to a fixed-length representation $\bm{f}_{v}$, and the text encoder $\mathcal{F}_{\theta_{t}}$, which processes text input, $T$, and generates a latent textual feature $\bm{f}_{t}$. In zero-shot T2I retrieval, a string of text, $q$, is used to query a pool of images, $\mathcal{I}$. Given the representations of the query, $\bm{f}_{q}$, and of all images $\{\bm{f}_{v,i}\}_{i\in\mathcal{I}}$, a cosine similarity score $s_{i}=\mathtt{sim}(\bm{f}_{q},\bm{f}_{v,i})$ is computed for each image, $i$, to obtain a final ranking.
@@ 3.2 EFSA: Episodic Few-Shot Adaptation
Although CLIP-style models achieve impressive zero-shot performance across varied downstream tasks, they struggle to accurately rank hard negative pairs in T2I retrieval, particularly in complex, multi-domain scenarios. We address this limitation with EFSA, an adaptation framework which utilizes top-$k$ highly similar images as hard negatives, to improve multi-modal alignment in the specific query domain. In this section, we detail our three-stage EFSA methodology, explaining how each stage contributes to more robust retrieval across diverse and challenging domains. For a visualization of the approach, see Figure 2.
Initial Top-$k$ Retrieval In the initial zero-shot retrieval step, we compute the score, $s$, between a given query text, q{q}, and all images $\bm{i}\in\mathcal{I}$. The top-$k$ most similar images are selected, forming the set $\mathcal{I}_{\text{top}}=\{\bm{i}_{1},\bm{i}_{2},\dots,\bm{i}_{k}\}$. Given a large-enough $K$ and a sufficiently generalized VLM, we can expect the ground truth to be in $\mathcal{I}_{\text{top}}$, but possibly not at the top rank, where it should be. Crucially, the rest of the images in the set would be similar to each other and to the ground truth, thus forming a set of domain-specific hard negatives with respect to the ground truth.
Data Augmentation and Episodic Training In this step, pseudo-captions $\mathcal{C}_{\text{top}}=\{\bm{c}_{1},\bm{c}_{2},\dots,\bm{c}_{k}\}$ are generated for each image in $\mathcal{I}_{\text{top}}$ using a pre-trained image captioning model, guided by a predefined prompt $\mathcal{P}$, chosen to ensure relevant caption generation. Together $\mathcal{I}_{\text{top}}$ and $\mathcal{C}_{\text{top}}$ form a dataset of image-caption pairs, which are used to adapt the VLM using two losses: contrastive and hinge.
@@
where $\text{sim}(\cdot,\cdot)$ denotes cosine similarity , and $\tau$ is the temperature parameter.
@@
The final test-time adaptation loss, $\mathcal{L}_{\text{test}}$, is a weighted combination of the two: $\mathcal{L}_{\text{test}}=\alpha\,\mathcal{L}_{\text{contrastive}}+\beta\,\mathcal{L}_{\text{hinge}}$ , where $\alpha$ and $\beta$ balance the contribution of each component.
These objectives are used to learn a set of LoRA layers [14] in both $\mathcal{F}_{\theta_{v}}$ and $\mathcal{F}_{\theta_{t}}$ in the VLM. With LoRA, the VLMs is adapted at a low computational cost, with reduced risk of overfitting to noise in $\mathcal{I}_{\text{top}}$ and $\mathcal{C}_{\text{top}}$.
Inference In this final step, the updated VLM is used to re-rank the images in $\mathcal{I}_{\text{top}}$ with respect to the query $q$, using cosine similarity and leveraging domain-adapted representations to enhance retrieval recall at the top ranks.
After each inference step, all LoRA weights are reset, ensuring that subsequent queries start from a neutral state, allowing the model to adapt to each new query independently, without cross-domain interference.
@@ 4.1 Experimental Setup
Implementation Details We use CLIP-B16 as the main VLM in all experiments, but additionally test the generalization of our approach to SigLIP [41] as well (see Supplementary §7). LoRA layers are added to all attention layers and multi-layer perceptron components, in both the vision and text encoders. The LoRA weights are initialized with Xavier initialization, using a rank of $r=64$ and a scaling factor of 15. The loss function is updated in a single step with parameters $\alpha=1.7$ and $\beta=0.3$, using the AdamW optimizer and a learning rate of 5e-4. All hyperparameters were tuned on the validation splits of COCO and Flickr30k, averaging scores over the two datasets.
For top-$k$ sampling, we set $k=16$. Pseudo-captions are generated with the LLaVA 1.5-13B model [22], using the prompt: “Describe what you see in detail with a maximum of 30 words.” An ablation over different prompts is reported in Supplementary §8. As the captions are independent of the queries, they can be generated offline and cached. Similarly, we precompute and cache the representations of all images in the retrieval pool, as every initial top-$k$ retrieval step relies on the same pre-trained VLM representations.
@@
Experimental Settings We study two experimental settings: a single-domain setting, where the pool contains only images from a single dataset; and a more challenging multi-domain setting, where all datasets in Table 1 are mixed into one large pool. The evaluation sets (e.g., COCO, Flickr30k) remain unchanged, but queries $\mathit{X}$ from domain $\mathit{Y}$ must retrieve their ground-truth among distractors from other domains $\mathit{Y^{\prime}}$, simulating realistic open-domain conditions. Images from datasets like CC12M serves solely as distractors, simulating realistic open-domain conditions where the system must retrieve relevant items from a large, diverse, and noisy collection. We recognize that enlarging the retrieval pool may introduce semantically relevant images not labeled as ground-truth (i.e., potential false negatives.) However, since all models in our study are evaluated under the same condition, comparisons remain fair. A robust retriever should be able to pick up even a weak signal, if present, and rank the true match higher than other semantically similar images.
@@
Below, we compare our approach to these baselines using the standard retrieval metric Recall@kk [^2] We use Recall@$k$ following prior work [7, 18, 19, 44]. For single ground truth, precision@$k$ is approximately equivalent to Recall@kk/$k$, and mAP is better suited for multi-target scenarios, which is not the case here. for $k=1,5,10$, with special attention to Recall@1, since EFSA focuses on an optimal prediction in the top-$1$ position.
@@ 4.2 Results
The main results are shown in Table 2 for the single-domain setting, and Table 3 for the multi-domain setting, in which the retrieval pool spans multiple diverse domains.
Single-Domain Setting While this setting is not the ultimate target of our work, it lays the foundation for the follow-up discussion of the multi-domain results. The first thing to note here is the considerable variation in zero-shot scores across the different datasets. The Books, ArtCap and SciCap datasets, in particular, prove far more challenging than COCO and Flickr30k, underscoring the importance of domain diversity in T2I retrieval evaluation.
@@
RLCF delivers improvements over the zero-shot baseline for 6 out of 8 datasets, but these improvements are mostly negligible in comparison to those seen with the fine-tuning baseline. The one exception beign SciCap where RLCF indeed proves best overall, yielding a Recall@1 improvement of 6.54 points. Meanwhile, however, for NASA and TextCap, RLCF exhibits diverging behavior, with scores plummeting for these datasets across all Recall@k metrics. This instability in the reinforcement-based optimizaiton of RLCF leaves simple fine-tuning as the strongest baseline on average for us to compare against.
@@
In the multi-domain setting, EFSA proves best on 6 out of 8 datasets, with an average Recall@1 4.27 points over the zero-shot baseline (27.56 v. 23.29). This number coincidentally matches exactly the improvement of EFSA over the zeros-shot baseline in the single-domain setting. In light of the absolute drop in scores in the multi-domain setting compared to the single-domain setting, this stable improvement with EFSA indicates that our method is highly robust to noise in the $\mathcal{I}_{top}$ subset: even if some lower-quality candidates get retrieved, EFSA successfully adapts to the domain-specific patterns and ranks better candidates higher. The same trend is observed when EFSA is applied to a SigLIP backbone (Supplementary Table 6).
@@ 4.3 Qualitative Analysis
Figure 3 provides an insight into the performance gains achieved with EFSA. Looking at the top-4 predictions retrieved with the zero-shot method and with EFSA for two queries from the Flickr30k dataset, we see that EFSA picks up on fine details such as the presence of a wine glass in the top image, and the orientation of the man in the bottom one.
@@
4.4 Computational & Storage Costs
Like most TTA methods, EFSA introduces additional computation compared to zero-shot retrieval. Yet, it remains more efficient than prior TTA methods such as RLCF. For example, on the COCO test set (5,000 queries) with $k=8$, EFSA takes 47 minutes, compared to 52 minutes for RLCF in the same setting. In the multi-domain setting, as the retrieval pool size increases, the computation scales linearly with the size of the pool, for both the zero-shot retrieval and the initial top-$k$ retrieval step in the TTA methods. In the subsequent adaptation step, however, the computational cost is independent of the pool size and is instead governed by the $k$ parameter. Since the synthetic captions are generated and cached in advance, there is no added latency from this step at inference time. Although EFSA stores additional image captions compared to the zero-shot baseline, the resulting storage overhead is negligible (see Appendix 10.)
Figure 3: Qualitative comparison of EFSA and zero-shot CLIP. Ground-truth images (highlighted in green) correspond to the text queries shown above each row. EFSA successfully re-ranks the ground-truth images to the first rank. See Supplementary Figure 7 for more detailed examples with generated captions.
@@ 5 Ablations
We perform extensive empirical analysis and ablation studies on COCO, NASA, SciCap, and ArtCap to assess the impact of key design choices, including fine-tuning parameters, top-$k$ selection, number of adaptation epochs, and the image captioner. An additional ablation on the loss function is provided in the supplementary material (Appendix 11.)
@@
Figure 4: Comparison of LoRA parameter tuning versus tuning all model parameters across multiple datasets
Table 4: Comparison of recall performance at incremental top-$k$ values across datasets.
@@ 5.2 Top-k Selection
The Recall@1 metric exhibits a more nuanced pattern: while the scores initially improve as the top-$k$ value increases, the gains plateau or even decline beyond a certain top-$k$. For example, on COCO, Recall@1 reaches a peak at $k=16$ with a score of 30.14 but drops slightly as $k$ continues to increase. This suggests that while expanding the retrieval pool improves general recall, very large top-$k$ values may introduce additional noise, compromising precision at the top rank. Importantly, since performance gains tend to saturate at moderate $k$ values, one can default to a lower $k$ (e.g., 8) in compute-constrained settings without significant performance degradation.
@@ 5.3 Effect of Epoch
All experiments so far were performed with a single epoch of training. In Figure 5 we explore whether increasing the number of epochs has a positive impact on performance, with the finding that by and large that is not the case. For two datasets, COCO and ArtCap, we see slightly higher scores at Recall@1 and Recall@5 on the second epoch of training, but the general trend is for recall to drop with more extensive training. Interestingly, the drop is more pronounced at higher ranks. In NASA, for example, Recall@1 drops by less than 10 points across the four epochs of training, while Recall@5 drops by over 25 points. This indicates that prolonged training on a limited or domain-specific set of images can cause the model to memorize specific features rather than develop more generalized representations, robust across various types of images. Regardless, from a computational point of view, having to perform a single epoch of training to reap the benefits of EFSA, is optimal.
5.4 Effect of Image Captioner
In Table 5 we measure EFSA’s performance with captions generated with LLaVA-13B, LLaVA-7B [22] and TinyLLaVA [45], and find that the choice of captioning model is not critical. Lighter models can be used to reduce computational overhead without a substantial change in results. To further assess the effectiveness of the synthetic captions, we compare their retrieval performance against ground-truth captions. Specifically, in the episodic fine-tuning we replace the synthetic captions with the ground-truth captions to obtain the results shown in Row 1 of Table 5. The minimal gain in performance confirms that synthetic captions effectively capture image semantics and serve as reliable substitutes for ground-truth captions.
Table 5: Retrieval performance using different captioning models.
Captions from COCO ArtCap
R@1 R@5 R@10 R@1 R@5 R@10
Ground-truth 40.64 64.64 72.61 19.92 38.29 45.31
LLaVA-13B 40.41 65.01 72.89 19.93 38.49 45.91
LLaVA-7B 40.33 64.96 72.69 19.95 38.40 45.71
TinyLLaVA-3.1B 39.47 64.39 72.79 19.38 37.56 45.18